Find Fourier integral of 𝑓(π‘₯) = 𝐻(π‘₯ βˆ’ 1) βˆ’ 𝐻(π‘₯ βˆ’ 2)


1
Expert's answer
2020-10-26T20:24:35-0400
"\dfrac{1}{\sqrt{2\pi}}\displaystyle\int_{-\infin}^\infin(H(t-1)-H(t-2))e^{iwt}dt="

"=\dfrac{1}{\sqrt{2\pi}}\delta(w)+\dfrac{ie^{iw}}{\sqrt{2\pi }w}-\big(\dfrac{1}{\sqrt{2\pi}}\delta(w)+\dfrac{ie^{2iw}}{\sqrt{2\pi }w}\big)="

"=\dfrac{ie^{iw}(1-ie^{iw})}{\sqrt{2\pi }w}"